In each of the 4 races, the democrats have 60% chance of winning.

Assuming that the races are independent of each other, what is the

probability that:

The Democrats will win a majority of the races.


1
Expert's answer
2021-12-09T14:00:03-0500

Let "X=" the number of wins of the Democrats out of 4 races:

X"\sim Bin(n,p)"

Given "n=4, p=0.60, q=1-p=0.40."


"P(X>2)=P(X=3)+P(X=4)"

"=\dbinom{4}{3}(0.6)^3(0.4)^{4-3}+\dbinom{4}{4}(0.6)^4(0.4)^{4-4}"

"=0.3456+0.1296=0.4752"